How much water must be added to 3 gallons of a 4% insecticide solution to reduce the concentration to 3%.
Answer by stanbon(75887)

How much water must be added to 3 gallons of a 4% insecticide solution to reduce the concentration to 3%.
--------
Equation:
active + active = active
0.04*4 + 0*x = 0.03(4+x)
---
Multiply thru by 100 to get:
4*4 + 0 = 3*4 + 3x
3x = 4
x = 4/3 gallons (amt of water needed to reduce the concentration).
-----------
Cheers,
Stan H.